2063 0.9 m 9000 N Aluminum Composite Panel (ACP) is a type of flat panel that consists of two thin aluminum sheets bonded to a non-aluminum core. ACPs are frequently used for external cladding of buildings (building facades), for insulation and for signage. Epcots Spaceship Earth is an example of the use of ACP in architecture. It is a geodesic sphere composed of 11,324 ACP tiles. A 0.9 m long by 9 mm thick panel is constructed of two 0.4 mm thick aluminum panels with a low density polyethylene (LDPE) core. A 9000 Newton load is applied to the panel. The modulus of elasticity of the aluminum is 69 GPa, while the modulus of elasticity of the LDPE is 0.2 GPa EXAMPLE NOTES IMAGES DISCUSS UNITS STATS HELP Part Description Answer Chk History What is the cross sectional area of the LDPE core? (include units with answer) 6.47 pts, 110% 2% try penalty A. # tries: 0 Show Details Format Check Hints: 1.0 What is the stress in the aluminum? (include units with answer) 6.47 pts. 1 10% # tries: 0 Show Details 8 Format Check-I 2% try penalty What is the stress in the LDPE core? (include units with answer) Format Check 6.47 pts. 110% ! 2% try penalty C. | # tries: 0 Show Details Foreck -1 296 try Previous Next
Answer #1

a) area = 9mm x 0.9m = 9 x 10-3 x 9 x 10-1 = 81 x 10-4 m2

c)stress LDPF = Fnormal / Area = 9000/(81x10-4) = 1.1111111 GPa

b)Stress Aluminium = strain x modulus of elasticity = 69 GPa x strain = 0.69 GPa (assuming strain = 1/100)

in case of any doubt, please do comment sir.